# BERT or FastText? A Comparative Analysis of Contextual as well as Non-Contextual Embeddings

### v2

@@ Abstract
Natural Language Processing (NLP) for low-resource languages presents significant challenges, particularly due to the scarcity of high-quality annotated data and linguistic resources. The choice of embeddings plays a critical role in enhancing the performance of NLP tasks, such as news classification, sentiment analysis, and hate speech detection, especially for low-resource languages like Marathi. In this study, we investigate the impact of various embedding techniques—Contextual BERT-based, Non-Contextual BERT-based, and FastText-based on NLP classification tasks specific to the Marathi language. Our research includes a thorough evaluation of both compressed and uncompressed embeddings, providing a comprehensive overview of how these embeddings perform across different scenarios. Specifically, we compare two BERT model embeddings, Muril and MahaBERT, as well as two FastText model embeddings, IndicFT and MahaFT. Our evaluation includes applying embeddings to a Multiple Logistic Regression (MLR) classifier for task performance assessment, as well as TSNE visualizations to observe the spatial distribution of these embeddings. The results demonstrate that contextual embeddings outperform non-contextual embeddings. Furthermore, BERT-based non-contextual embeddings extracted from the first BERT embedding layer yield better results than FastText-based embeddings, suggesting a potential alternative to FastText embeddings.
BERT or FastText? A Comparative Analysis of Contextual as well as Non-Contextual Embeddings
@@ 1 Introduction
Word embedding is a way of representing words into dense vectors in a continuous space such that the vectors capture the semantic relationship between the words for the models to understand the context and meaning of the text. FastText, a context-independent method, basically captures the subword information, enabling it to learn rare words, misspelled words as well as out-of-vocabulary words. It is recognized in the NLP community for its efficient performance in tasks like text classification and sentiment analysis. Despite being relatively old, it still remains one of the most effective alternatives when performing tasks on large datasets across various languages due to its subword-based approach.
@@
BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018) word embeddings understand the meaning of a word based on its context in a sentence. This has led to improved performance for various tasks. Numerous variations of BERT like IndicBERT Kakwani et al. (2020), MuRIL Khanuja et al. (2021), AfriBERT Ralethe (2020), and mBERT Devlin et al. (2018) to name a few, are available for experiments.
@@
Recent studies have experimented with both FastText and BERT for various tasks. Experiments of D’Sa et al. (2020) demonstrated that BERT embeddings outperformed FastText for classifying English text into toxic and non-toxic. Findings of Ahmed et al. (2024) suggested that BERT embeddings outperformed those of FastText with an F1 score of 84% when evaluated for depressive post-detection in Bangla.
@@
While BERT consistently outperforms other word embeddings in various tasks in high-resource languages (HRLs) like English, its efficacy on low-resource languages (LRLs) remains relatively underexplored. Specifically, the comparison of BERT and FastText embeddings for tasks in LRLs like news classification, sentiment analysis, and hate speech detection still warrants further investigation to assess BERT’s effectiveness across a wider range of languages and tasks. This gap is particularly pronounced when balancing model performance with computational efficiency, which becomes a critical factor in low-resource settings.
Due to limited work in low-resource languages, it is not well studied as to which one will perform better in LRL scenarios. While BERT contextual embeddings are shown to work better than non-contextual FastText, it has significant computational overhead as compared to the FastText model. In our work, we focus on investigating how well non-contextual BERT embeddings from the first embedding layer of the model perform in comparison to FastText. Further, since BERT embeddings are 768-dimensional while FastText embeddings are 300-dimensional, we perform Singular Value Decomposition (SVD)-based compression on BERT embeddings to examine the performance of compressed BERT embeddings.
@@
This paper focuses on utilizing FastText and BERT for the Marathi language for the following tasks: Sentiment Classification, 2-Class and 4-Class Hate Speech Detection, and News Article Classification for headlines, long paragraphs, and long documents. We construct a comprehensive analysis of FastText embeddings, IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings, and BERT embeddings, including muril-base-cased Khanuja et al. (2021) and marathi-bert-v2 Joshi (2022a). Additionally, we experimented on compressed as well as contextual and non-contextual variants of the BERT-based model embeddings. This analysis revealed that contextual BERT embeddings consistently outperformed FastText across all evaluated tasks. Additionally, non-contextual BERT embeddings demonstrated superior performance to FastText in most tasks.
@@
Conducting a comprehensive study on BERT and FastText embeddings for Marathi, a low-resource language, across diverse classification tasks, including sentiment analysis, news classification, and hate speech detection.
Evaluating the impact of embedding compression by comparing the performance of compressed and uncompressed embeddings.
Investigating the differences between contextualized and non-contextualized representations in BERT embeddings.
Examining the trend between non-contextual BERT and FastText embeddings.
The paper is organised as follows: Section 2 provides a concise review of prior research on FastText and BERT. Section 3 includes the datasets and model embeddings that are utilized for the experiments. Section 4 presents the methodology used. Section 5 presents the results and key insights drawn from the findings along with a comparative analysis of FastText embeddings and BERT. In Section 6, we conclude our discussion.
@@ 2 Literature Review
The existing literature emphasizes the superiority of BERT embeddings over traditional word embedding techniques like Word2Vec Mikolov et al. (2013) , GloVe Pennington et al. (2014), and FastText across various natural language processing (NLP) tasks. For instance, Khaled et al. (2023) compare four popular pre-trained word embeddings—Word2Vec (via Aravec Mohammad et al. (2017)), GloVe, FastText, and BERT (via ARBERTv2)—on Arabic news datasets. They highlight BERT’s superior performance, achieving over 95% accuracy due to its contextual interpretation.
@@
Similarly, Kabullar and Türker (2022) analyzes the performance of embeddings on the AG News dataset, which includes 120K instances across four classes. They conclude that BERT outperforms other methods, achieving 90.88% accuracy. FastText, Skip-Gram, CBOW, and GloVe achieve 86.91%, 85.82%, 86.15%, and 80.86%, respectively.
@@
While traditional embeddings perform reasonably well, the consistent dominance of BERT in complex tasks is also noted in sentiment analysis. For instance, Xie et al. (2024) explores how combining BERT and FastText embeddings enhances sentiment analysis in education, demonstrating that BERT’s contextual understanding, along with FastText’s ability to handle out-of-vocabulary words, improves generalization over unseen text.
@@
In the domain of toxic speech classification, D’Sa et al. (2020) utilize both BERT and FastText embeddings to classify toxic comments in English, with BERT embeddings outperforming FastText. This trend continues in hate speech detection, where Rajput et al. (2021) find that neural network classifiers using BERT embeddings perform better than those with FastText embeddings alone, further supporting BERT’s effectiveness.
@@
Additionally, Chanda (2021) assess BERT embeddings against traditional context-free methods (GloVe, Skip-Gram, and FastText) for disaster prediction, demonstrating BERT’s superior performance in combination with traditional machine learning and deep learning methods.
@@
For low-resource languages (LRLs), Ahmed et al. (2024) examine methods like traditional TF-IDF, BERT, and FastText embeddings within a CNN-BiLSTM architecture for detecting depressive texts in Bangla. Their results show that BERT embeddings yield the highest F1 score (84%), indicating their dominance over other methods. This suggests that BERT’s efficacy extends even to LRLs.
@@
In medical applications, Khan et al. (2024) proposes integrating BERT embeddings with SVM for prostate cancer prediction. By incorporating both numerical data and contextual information from clinical text, they achieve 95% accuracy, far outperforming the 86% accuracy achieved with numerical data alone.
@@
Moreover, Malik et al. (2021) uses both BERT and FastText embeddings to preprocess a dataset of conversations from Twitter and Facebook. Applying various machine learning and deep learning algorithms, they find that CNN yields the best results, further demonstrating BERT’s capabilities.
@@
Finally, while Asudani et al. (2023) offers a comprehensive analysis of traditional word embeddings alongside more advanced techniques like ELMo and BERT, providing insights into commonly used datasets and models for benchmarking, Umer et al. (2022) highlight FastText’s versatility in various domains, despite BERT’s consistently superior performance.
We note that the reviewed literature underscores the consistent superiority of BERT embeddings across diverse NLP tasks and domains. However, most existing work predominantly focuses on high-resource languages, emphasizing the need for further exploration and adaptation of these techniques for low-resource languages to achieve advancements in NLP.
@@ 3 Datasets and Models Used
In our research work, we used 3 Marathi datasets MahaSent: A 3-class sentiment analysis dataset Pingle et al. (2023), MahaHate: Includes both a 2-class as well as a 4-class hate classification dataset Patil et al. (2022) and MahaNews: A news categorization dataset with 12 classes Mittal et al. (2023).
@@ 4 Methodology
For each sentence, corresponding embeddings were generated. The creation of BERT embeddings was done by first tokenizing the text using the BERT tokenizer, along with padding and truncation. The tokenized input was then passed to the model and the output of the last hidden layer of BERT was taken, which was then averaged to get contextual embeddings for every sentence. Whereas for non-contextual embeddings, the output of the first embedding layer was used. Refer Fig 1 for the embedding extraction workflow diagram.
@@
Singular Value Decomposition (SVD) to select the most relevant features, extracting the top 300 components for all the combinations of contextual as well as non-contextual for MahaBERT as well as Muril. Refer Fig 2 for compression flow diagram. All embeddings were then passed to a multiple logistic regression(MLR) classifier for classification into target labels.
To determine the validity of the results obtained, 5-fold cross-validation was performed for all tasks, and the results are presented in Table 2.
@@ 5 Results
In sections 5.1 and 5.2, we have considered the uncompressed versions of Muril and MahaBERT. Further, in section 5.3, we specifically show the effect of compression on Muril and MahaBERT. In Section 5.4, we identify patterns in the embedding visualizations shown in Figure 3.
@@ 5.3 Effect of Compression
From Table 1, we can also infer that compression has a negative effect on non-contextual embeddings, as uncompressed ones tend to perform better. However, the impact of compression on contextual embeddings varies across datasets, making it difficult to draw a consistent conclusion.
5.4 Inference from TSNE
Analyzing the t-SNE embedding visualizations for LDC specifically, BERT-based embeddings (MahaBERT and Muril) tend to form more compact clusters compared to FastText-based embeddings. Although some class overlap exists, BERT embeddings show denser class formations.
@@ 6 Conclusion
In our research, we analyzed the effectiveness of various BERT and FastText-based embeddings on three key NLP tasks for Marathi: news classification, hate speech classification, and sentiment classification. Additionally, we examined the spatial distributions of embeddings generated for one of the classification task.
Our study shows that contextual MahaBERT embeddings generally outperform contextual Muril as well as FastText-based embeddings across the 3 tasks. Our results indicate that contextual embeddings yield better outcomes than non-contextual ones .
Regarding the effect of compression: compressing non-contextual embeddings has a negative impact on performance, whereas compressing contextual embeddings shows no definitive effect.
When comparing non-contextual BERT embeddings to FastText embeddings, the non-contextual BERT embeddings outperform the FastText embeddings in the majority of tasks.
7 Acknowledgement

### v3

# Non-Contextual BERT or FastText? A Comparative Analysis

@@ Abstract
Natural Language Processing (NLP) for low-resource languages, which lack large annotated datasets, faces significant challenges due to limited high-quality data and linguistic resources. The selection of embeddings plays a critical role in achieving strong performance in NLP tasks. While contextual BERT embeddings require a full forward pass, non-contextual BERT embeddings rely only on table lookup. Existing research has primarily focused on contextual BERT embeddings, leaving non-contextual embeddings largely unexplored. In this study, we analyze the effectiveness of non-contextual embeddings from BERT models (MuRIL and MahaBERT) and FastText models (IndicFT and MahaFT) for tasks such as news classification, sentiment analysis, and hate speech detection in one such low-resource language—Marathi. We compare these embeddings with their contextual and compressed variants. Our findings indicate that non-contextual BERT embeddings extracted from the model’s first embedding layer outperform FastText embeddings, presenting a promising alternative for low-resource NLP.
Non-Contextual BERT or FastText? A Comparative Analysis
@@ 1 Introduction
Word embedding is a way of representing words into dense vectors in a continuous space such that the vectors capture the semantic relationship between the words for the models to understand the context and meaning of the text. FastText, a context-independent method, basically captures the subword information, enabling it to learn rare words, misspelled words, and out-of-vocabulary words. It is recognized in the NLP community for its efficient performance in tasks like text classification and sentiment analysis. Despite being relatively old, it still remains one of the most effective alternatives when performing tasks on large datasets across various languages due to its subword-based approach.
@@
BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2018) word embeddings understand the meaning of a word based on its context in a sentence. The embeddings extracted just before the first embedding layer of the BERT architecture are referred to as non-contextual embeddings, while those obtained from the last hidden layer of BERT are known as contextual embeddings (Refer Fig 1). Numerous variations of BERT like IndicBERT Kakwani et al. (2020), MuRIL Khanuja et al. (2021), AfriBERT Ralethe (2020), and mBERT Devlin et al. (2018) to name a few, are available for experiments.
@@
Recent studies have experimented with both FastText and BERT for various tasks; however, most of them focus on exploring contextual BERT embeddings. Experiments of D’Sa et al. (2020) demonstrated that BERT embeddings outperformed FastText for classifying English text into toxic and non-toxic. Findings of Ahmed et al. (2024) suggested that BERT embeddings outperformed those of FastText with an F1 score of 84% when evaluated for depressive post-detection in Bangla.
@@
While BERT consistently outperforms other word embeddings in various tasks for high-resource languages (HRLs) like English (Malik et al. (2021)), its effectiveness in low-resource languages (LRLs) remains relatively underexplored. This gap is particularly pronounced when balancing model performance with computational efficiency, which becomes a critical factor in low-resource settings.
Previous studies (D’Sa et al. (2020)) have focused on contextual BERT embeddings, which outperform FastText due to their ability to capture contextual information. However, the use of non-contextual BERT embeddings for classification tasks in low-resource languages like Marathi remains unexplored. Unlike contextual embeddings, which require a full forward pass through the model, non-contextual embeddings can be obtained through a simple table lookup. To our knowledge, no prior work has examined the effectiveness of non-contextual BERT embeddings. We investigate how these embeddings, extracted from the model’s first layer, compare to FastText embeddings for tasks such as news classification, sentiment analysis, and hate speech detection in Marathi.
Additionally, past comparisons often used BERT’s 768-dimensional embeddings against FastText’s 300-dimensional ones, which is unfair since higher dimensions naturally provide better feature extraction. To address this, we ensure a fair comparison by reducing the BERT embeddings to 300 dimensions.
@@
This paper focuses on utilizing FastText and non-contextual BERT for the Marathi language for the following tasks: Sentiment Classification, 2-Class and 4-Class Hate Speech Detection, and News Article Classification for headlines, long paragraphs, and long documents. We construct a comprehensive analysis of FastText embeddings, IndicFT Kakwani et al. (2020) and MahaFT Joshi (2022a) embeddings, and BERT embeddings, including muril-base-cased Khanuja et al. (2021) and marathi-bert-v2 Joshi (2022a). To enhance the comparison, we replicate the experiments using widely utilized contextual BERT embeddings. We also evaluate the impact of compression on both contextual and non-contextual BERT-based embeddings. Our analysis shows that non-contextual BERT embeddings generally perform better than FastText in most tasks. Furthermore, contextual BERT embeddings consistently outperform FastText across all evaluated tasks. However, compressing non-contextual embeddings reduces their performance, making FastText more effective than compressed non-contextual BERT.
@@
We conduct a detailed study comparing non-contextual BERT embeddings and FastText embeddings for Marathi, a low-resource language. The evaluation covers multiple classification tasks, including sentiment analysis, news classification, and hate speech detection.
To ensure a fair comparison, we compress BERT embeddings from 768 to 300 dimensions using Singular Value Decomposition (SVD). This allows us to analyze how dimensionality reduction impacts BERT’s performance compared to its uncompressed version and FastText.
We explore the differences between contextual and non-contextual BERT embeddings, examining their impact on classification performance in low-resource settings.
The paper is organized as follows: Section 2 provides a concise review of prior research on FastText and BERT. Section 3 includes the datasets and model embeddings that are utilized for the experiments. Section 4 presents the methodology used. Section 5 presents the results and key insights drawn from the findings along with a comparative analysis of FastText embeddings and BERT. In Section 6, we analyze our results and explain the reasons behind them. In Section 7, we conclude our discussion.
@@ 2 Literature Review
The existing literature emphasizes the superiority of contextual BERT embeddings over traditional word embedding techniques like Word2Vec Mikolov et al. (2013) , GloVe Pennington et al. (2014), and FastText across various natural language processing (NLP) tasks. For instance, Khaled et al. (2023) compare four popular pre-trained word embeddings—Word2Vec (via Aravec Mohammad et al. (2017)), GloVe, FastText, and contextual BERT (via ARBERTv2)—on Arabic news datasets. They highlight BERT’s superior performance, achieving over 95% accuracy due to its contextual interpretation.
@@
Similarly, Kabullar and Türker (2022) analyzes the performance of embeddings on the AG News dataset, which includes 120K instances across four classes. They conclude that contextual BERT outperforms other methods, achieving 90.88% accuracy. FastText, Skip-Gram, CBOW, and GloVe achieve 86.91%, 85.82%, 86.15%, and 80.86%, respectively.
@@
While traditional embeddings perform reasonably well, the consistent dominance of contextual BERT in complex tasks is also noted in sentiment analysis. For instance, Xie et al. (2024) explores how combining BERT and FastText embeddings enhances sentiment analysis in education, demonstrating that BERT’s contextual understanding, along with FastText’s ability to handle out-of-vocabulary words, improves generalization over unseen text.
@@
In the domain of toxic speech classification, D’Sa et al. (2020) utilize both contextual BERT and FastText embeddings to classify toxic comments in English, with BERT embeddings outperforming FastText. This trend continues in hate speech detection, where Rajput et al. (2021) find that neural network classifiers using contextual BERT embeddings perform better than those with FastText embeddings alone, further supporting BERT’s effectiveness.
@@
Additionally, Chanda (2021) assess contextual BERT embeddings against traditional context-free methods (GloVe, Skip-Gram, and FastText) for disaster prediction, demonstrating BERT’s superior performance in combination with traditional machine learning and deep learning methods.
@@
For low-resource languages (LRLs), Ahmed et al. (2024) examine methods like traditional TF-IDF, contextual BERT, and FastText embeddings within a CNN-BiLSTM architecture for detecting depressive texts in Bangla. Their results show that BERT embeddings yield the highest F1 score (84%), indicating their dominance over other methods. This suggests that BERT’s efficacy extends even to LRLs.
@@
In medical applications, Khan et al. (2024) proposes integrating contextual BERT embeddings with SVM for prostate cancer prediction. By incorporating both numerical data and contextual information from clinical text, they achieve 95% accuracy, far outperforming the 86% accuracy achieved with numerical data alone.
@@
Moreover, Malik et al. (2021) uses both contextual BERT and FastText embeddings to preprocess a dataset of conversations from Twitter and Facebook. Applying various machine learning and deep learning algorithms, they find that CNN yields the best results, further demonstrating BERT’s capabilities.
@@
Finally, while Asudani et al. (2023) offers a comprehensive analysis of traditional word embeddings alongside more advanced techniques like ELMo and contextual BERT, providing insight into commonly used datasets and models for benchmarking, Umer et al. (2022) highlights the versatility of FastText in various domains, despite BERT’s consistently superior performance.
We note that the reviewed literature highlights the consistent superiority of BERT embeddings across various NLP tasks and domains. However, most existing studies focus mainly on contextual BERT embeddings, but not on non-contextual embeddings. Moreover, these studies predominantly address high-resource languages, leaving low-resource languages like Marathi largely unexplored. In particular, there is a lack of research assessing the effectiveness of non-contextual BERT embeddings for Marathi. Additionally, the impact of dimensionality leveling, i.e. the efficacy of BERT embedding compression, has not been explored.
@@ 3 Datasets and Models Used
In our research work, we used 3 Marathi datasets, MahaSent: A 3-class sentiment analysis dataset Pingle et al. (2023), MahaHate: A 2-class as well as a 4-class hate classification dataset Patil et al. (2022) and MahaNews: A news categorization dataset with 12 classes Mittal et al. (2023).
@@ 4 Methodology
For each sentence, corresponding embeddings were generated and the corresponding categorical labels were encoded into numerical labels. The creation of BERT embeddings was done by first tokenizing the text using the BERT tokenizer, along with padding and truncation. The tokenized input was then passed to the model and the output of the last hidden layer of BERT was taken, which was then averaged to get contextual embeddings for every sentence. Whereas for non-contextual embeddings, the output of the first embedding layer was used. Refer Fig 1 for the embedding extraction workflow diagram.
@@
Singular Value Decomposition (SVD) to select the most relevant features, extracting the top 300 components for all the combinations of contextual as well as non-contextual for MahaBERT as well as Muril. Refer Fig 2 for compression flow diagram.
Feature scaling was also applied to the outputs. All embeddings were then passed to a multiple logistic regression(MLR) classifier for classification into target labels.
4.1 Experimental Setup
The experiments were conducted on Kaggle notebooks equipped with a P100 GPU accelerator, utilizing 16 GB of GPU memory, 20 GB of storage, and 32 GB of RAM. Accuracy was chosen as the evaluation metric, given the balanced nature of the datasets. For classification, the results obtained from the embeddings were mapped to final labels using a multinomial logistic regression model to maintain methodological simplicity. To determine the validity of the results obtained, 5-fold cross-validation was performed for all tasks, and the results are presented in Table 2.
@@ 5 Results
In sections 5.1 and 5.2, we have considered the uncompressed versions of Muril and MahaBERT. Further, in section 5.3, we specifically show the effect of compression on Muril and MahaBERT.
@@ 5.3 Effect of Compression
From Table 1, it can be inferred that compression negatively impacts non-contextual embeddings, as uncompressed versions generally perform better. This is evident from MahaFT outperforming the compressed non-contextual MahaBERT embeddings in all datasets except MahaHate-4c, suggesting that compression lowers the performance of non-contextual BERT embeddings.
However, the effect of compression on contextual embeddings varies across datasets, making it challenging to derive a consistent conclusion.
6 Inference
In this section, we explain why the non-contextual MahaBERT embeddings outperform FastText (MahaFT and IndicFT) embeddings. Both MahaBERT and MahaFT embeddings have been trained on the same corpus of 752 million tokens Joshi (2022a). The superior performance of non-contextual MahaBERT embeddings can be attributed to its larger embedding size, training data size, and contextual training objective. Specifically, the embedding size for Marathi-BERT-v2 is 152M (197,285 × 768), compared to MahaFT, which is 132M (439,247 × 300).
IndicFT performs worse than MahaFT, likely due to its smaller dataset size of 551 million tokens Kakwani et al. (2020). On the other hand, contextual BERT achieves better results because its hidden layers are effectively utilized.
Additionally, we observe a negative impact when compressing MahaBERT non-contextual embeddings. Reducing the embedding size from 152M (197,285 × 768) to 59M (197,285 × 300) leads to a decrease in performance, likely due to the loss of representational capacity.
@@ 7 Conclusion
In our research, we analyzed the effectiveness of various BERT and FastText-based embeddings on three key NLP tasks for Marathi: news classification, hate speech classification, and sentiment classification focusing primarily on non-contextualised BERT embeddings.
Our results show that contextual BERT embeddings perform better than non-contextual ones, including both non-contextual BERT embeddings and FastText. Among non-contextual embeddings, BERT generally outperforms FastText in most tasks. However, when non-contextual BERT embeddings are compressed, their performance drops, and FastText performs better than compressed non-contextual BERT.
Acknowledgement
@@
Appendix A Appendix
A.1 T-SNE Visualisations
Refer to Fig. 4 for the t-SNE visualization of all datasets.